OpenStudy (anonymous):

What are two numbers for which the sum of there square is 628 and the differance of there squares is 340?

OpenStudy (anonymous):

x^2+y^2=628 x^2-y^2=360

OpenStudy (anonymous):

solce to get the answer

OpenStudy (saifoo.khan):

a^2 + b^2 = 628 a^-b^2 = 340

OpenStudy (anonymous):

let the numbers be a,b given \[a ^{2}+b ^{2}=628\]\[a ^{2}-b ^{2}=340\] then 2a^2=968 a^2=484 a=22 b=12

OpenStudy (anonymous):

damnit Srir beat me :(

OpenStudy (anonymous):

who?me?

OpenStudy (anonymous):

yeah you answered faster :P How did you know to add 340 to 628?

OpenStudy (anonymous):

na.....its not the matter of speed its the matter of accuracy....see i'm not smart than anyone...so u too won!